# Diophantine Approximation with Piatetski-Shapiro Primes

### v1

@@ Abstract.
We prove that for every irrational number $\alpha$, real number $c$ satisfying $81<c<9/$ 8 and positive real number $\theta$ satisfying $\theta<(9/c-8)/10$, there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p||<p^{-\theta}$.
@@ 1. Introduction
Fix an irrational number $\alpha$, a real number $c$ satisfying $111<c<12/$ 11 and a positive real number $\theta$ satisfying $\theta<(12/c-11)/26$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p||<p^{-\theta}$.
@@
Fix an irrational number $\alpha$, a real number $c$ satisfying $81<c<9/$ 8 and a positive real number $\theta$ satisfying $\theta<(9/c-8)/10$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p||<p^{-\theta}$.
@@ 2. Simplification and application of Harman’s sieve
$\sharp\left\{p\in\mathbb{P}:X/2\leq p<X,\ p=\left[n^{c}\right]\mbox{ for some }n\in\mathbb{N},\ ||\alpha p||<\Delta\right\}>0$
@@
$\mathcal{A}:=\left\{a\in\mathbb{N}:X/2\leq a<X,\ ||\alpha a||<\Delta,\ ||a^{\gamma}+2\delta||<\delta\right\}.$ (2.7)
@@ 3. Fourier analysis
We shall detect the conditions $||\alpha a||<\Delta$ and $||a^{\gamma}+2\delta||<\delta$ in (2.7) using the following Fourier analytic device.
@@
$\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}\Xi^{-}\left(\alpha mn,(mn)^{\gamma}+2\delta\right)\leq\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ m\leq X^{15/22}\end{subarray}}a_{m}\leq\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}\Xi^{+}\left(\alpha mn,(mn)^{\gamma}+2\delta\right)$
@@
$\begin{split}\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\Xi^{-}\left(\alpha mn,(mn)^{\gamma}+2\delta\right)\leq&\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\\ \leq&\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\Xi^{+}\left(\alpha mn,(mn)^{\gamma}+2\delta\right).\end{split}$

### v2

@@ Abstract.
We prove that for every irrational number $\alpha$, real number $\beta$, real number $c$ satisfying $81<c<9/$ 8 and positive real number $\theta$ satisfying $\theta<(9/c-8)/10$, there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
@@ 1. Introduction
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $111<c<12/$ 11 and a positive real number $\theta$ satisfying $\theta<(12/c-11)/26$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
@@
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $81<c<9/$ 8 and a positive real number $\theta$ satisfying $\theta<(9/c-8)/10$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
@@ 2. Simplification and application of Harman’s sieve
$\sharp\left\{p\in\mathbb{P}:X/2\leq p<X,\ p=\left[n^{c}\right]\mbox{ for some }n\in\mathbb{N},\ ||\alpha p+\beta||<\Delta\right\}>0$
@@
$\mathcal{A}:=\left\{a\in\mathbb{N}:X/2\leq a<X,\ ||\alpha a+\beta||<\Delta,\ ||a^{\gamma}+2\delta||<\delta\right\}.$ (2.7)
@@ 3. Fourier analysis
We shall detect the conditions $||\alpha a+\beta||<\Delta$ and $||a^{\gamma}+2\delta||<\delta$ in (2.7) using the following Fourier analytic device.
@@
$\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}\Xi^{-}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right)\leq\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ m\leq X^{15/22}\end{subarray}}a_{m}\leq\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}\Xi^{+}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right)$
@@
$\begin{split}\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\Xi^{-}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right)\leq&\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\\ \leq&\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\Xi^{+}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right).\end{split}$